new to this completely.-3/4t=-7/8?
Answer by checkley71(8403)

-3/4T=-7/8 DIVIDE BOTH SIDES BY -3/4
T=(-7/8)/(-3/4) NOW INVERT THE DENOMINATOR & MULTIPLY THESE FRACTIONS.
T=(-7/8*-4/3)
T=28/24
T=7/6 ANSWER.